An automobile manufacturer stocks a certain inventory item. The daily demand for the inventory item is 20 units. The company estimates its annual holding cost for this item to be $10 per unit. The cost to place and process an order from the supplier is $200. The company operates 300 days per year, and the lead time to receive an order from the supplier is 3 working days on average.

(a)    Find the economic order quantity.

(b)    Find the average and maximum inventory.

(c)    Find the total inventory cost.

(d)    What is the reorder point? What is the main limitation of using this reorder point?

(e)     Suggest an alternate inventory management system/model to the manufacturer indicating associated benefits.
Answer #1

Given that,

[ Ch] Holding Cost = $10 per unit

[Co ]Ordering Cost = $200 per order

[ D ]Annual Demand =300*20=6000

Formula for Economic Order Quantity =Square Root of [( 2 X D X Co)/ Ch],

Or, Economic Order Quantity = Square Root of [(2 X 6000 X 200)/ (10)]

Or, Economic Order Quantity = Square Root of [ 240000] = 489.898

Number of Order = Annual Demand / Optimal Order Quantity

Number of Order =6000/489.898=12.25

Total Annual Ordering Cost = Number of Order X Ordering Cost per Order = 12.25 X USD 200= $2449.49

Total Annual Holding Cost = Average Units X Holding Cost per unit per Year =(489.898/2)*10=2449.49

Total Inventory cost = Total Annual Ordering Cost+ Total Annual Holding Cost=$2449.49+$2449.49=$4898.98

Reorder Point=(Average daily demand X Given Lead Time) + Safety Stock=20*3+ Safety Stock=60 +Safety Stock

Without considering Safety Stock, Reorder point=60

1)

Economic Order Quantity = 490(Rounded)

2)

Average Inventory=490/2=245

Minimum Inventory = Reorder level-(Average Inventory*Lead time)=60-(245*3)=-675

Minimum inventory can’t be negative, hence, it is Zero

Maximum inventory= Economic Order Quantity + Reorder Quantity-(Minimum Inventory* Lead time) =490+60-0=550

3)

Total Inventory cost=$4898.98=$4899

4)

Without considering Safety Stock, Reorder point=60

5)

The Main Limitation of Reorder Point is if demand is much variable, then we could face difficult situation to estimate reorder point value. Reorder Point value should add safety stock, as without such safety stock, organization could be in trouble and face lack of inventory for their use. System could be overloaded if multiple items are reordered at the same time.

6)

Alternative Inventory management system or model to the manufacturer is Just in Time or JIT approach with this technique, as JIT system will work on pull base system, where, customer order gets defined first, then production and other operations are started knowing the customer orders.